A company makes electronic gadgets. One out of every 50 gadgets is faulty, but the company doesn't know which ones are faulty until a buyer complains. Suppose the company makes a Php 1,000.00 profit on the sale of any working gadget, but suffers a loss of Php 25, 000.00 for every faulty gadget because they have to repair the unit.


1
Expert's answer
2021-10-04T19:31:15-0400

There is Php 1,000.00 profit and Php 25, 000.00 loss on faulty gadgets.

We want to calculate if the company can expect profit in long run.

The probability of getting a faulty gadget is as follows.

P(Faulty)=1/50=0.02


P(Not Faulty)=1−0.02=0.98


The expected profit is as E(X)=∑xP(X=x).


E(X)=1000×0.98 +(−25000)×0.02=980-500=480


Thus, in a long run the company will have profit.